The local diner offers a meal combination consisting of an appetizer, a soup, a main course, and a dessert. There are six appetizers, four soups, six main courses, and six desserts. Your diet restricts you to choosing between a dessert and an appetizer. (You cannot have both.) Given this restriction, how many three-course meals are possible?
Answer #1

Choosing Appetizers, number of possible means = 6 * 4 * 6 = 144

Choosing Dessert, number of possible means = 4 * 6 * 6 = 144

Hence,

Total three course meals possible = 144 + 144 = 288